OpenStudy (anonymous):

PLEASE EXPLAIN: In this equation the variable z has two possible value is -2. What is the other possible value of z? -4|4z+2|=-24

OpenStudy (wikiemol):

You can solve this problem by setting 4z+2 = -6 and 4z+2 = 6

OpenStudy (anonymous):

yes, technically, to get rid of a "modulus", you have to square it

OpenStudy (anonymous):

\[ x^2=y\\ \implies |x|=\sqrt{y} \]

OpenStudy (anonymous):

@wikiemol what would be the next step. To subtract two from both sides?

OpenStudy (anonymous):

so the other answer to z would be 1? correct? @wikiemol

OpenStudy (wikiemol):

for the first equation: 4z=-6-2 4z=-8 z=-2 for the second equation 4z=6-2 4z=4 z=1